OpenStudy (inkyvoyd):

Isn't base ten a terrible base to do math in? I mean, it slows down computers, it's totally useless for expressing most rational functions, and, it's really freaking hard to convert!

OpenStudy (inkyvoyd):

Why 5 fingers? WHY?!

OpenStudy (anonymous):

Would you rather use hex? :D

OpenStudy (ash2326):

Real world works in base 10 so it's necessary to include in Computers. Even if they get slowed by that. We need Computers for our benefit , no matter they get slow :P

OpenStudy (inkyvoyd):

The real world doesn't use base ten. Just our minds.

OpenStudy (inkyvoyd):

I imagine that humans would have an easier time with everything if we used base 12 or base 16.

OpenStudy (inkyvoyd):

Or, we could just be babylonians.

hero (hero):

Base 60 is the best

OpenStudy (kinggeorge):

why not base 20? That was popular in the Americas.

OpenStudy (kinggeorge):

Fingers and toes!

OpenStudy (inkyvoyd):

base 20=base 2*2*5. There are too few rational numbers you can easily express.

OpenStudy (inkyvoyd):

Let's look at base 12: 1/2 in base ten=0.6 1/3 in base ten=0.4 1/4 in base ten=0.3 THe only disadvantage of base 12 is that you can't write 1/5 easily, but 1/5 is only important to us because we use base ten.

OpenStudy (kinggeorge):

I'm pretty sure I read some article suggesting that base e was the most efficient base to use. If I find the article I'll post a link.

OpenStudy (inkyvoyd):

Really?

OpenStudy (inkyvoyd):

does it have to do with e^(ipi)?

OpenStudy (inkyvoyd):

e is transcendental, so it might be good for pure math? but not for counting 0.o

OpenStudy (inkyvoyd):

"I have the natural logarithm of 1 dollar" (in base e)

OpenStudy (kinggeorge):

I was also thinking it sounded ridiculous for counting, but they seemed to be suggesting that. They made some effort to show it was the most efficient base to express larger numbers with. Or something like that. I read it a while back. :/

OpenStudy (inkyvoyd):

Or, I have (-1^(ipi)) dollars! MUAHAHHA

OpenStudy (kinggeorge):

So we should have about 2.7182818284590452353602874713526624977572470936999595... symbols to express numbers.

OpenStudy (inkyvoyd):

Hm.

OpenStudy (inkyvoyd):

I don't understand though, it's the most "optimal base", but it really sucks for counting.

OpenStudy (inkyvoyd):

obviously this guy didn't define optimal base very well.

OpenStudy (kinggeorge):

Interesting in theory, but absolutely terrible for application.

OpenStudy (inkyvoyd):

Yes. I bet it would be good for pure mathematics, but the number "1" becomes transcendental in base e!

OpenStudy (kinggeorge):

Shouldn't every rational number become transcendental? In fact, wouldn't every algebraic number become transcendental?

hero (hero):

Only God could use it

OpenStudy (inkyvoyd):

Exactly, it would be terrible.

OpenStudy (inkyvoyd):

And many common transcendental numbers would still be transcendental, I'm sure.

OpenStudy (inkyvoyd):

Like pi.

OpenStudy (kinggeorge):

I'm not actually sure if pi would remain transcendental or not. It's related to e from Euler's identity.

OpenStudy (inkyvoyd):

PI would remain transcendental.

OpenStudy (inkyvoyd):

e^(ipi) was used to prove that pi was transcendental

OpenStudy (inkyvoyd):

so, I'm pretty sure it would still be transcendental.

OpenStudy (kinggeorge):

But if e were algebraic, it would follow from that formula that pi was algebraic.

OpenStudy (inkyvoyd):

wait, we can't prove this question.

OpenStudy (inkyvoyd):

what we are looking for is essentially whether pi-e is transcendental.

OpenStudy (inkyvoyd):

I remember reading on wikipedia that it's not known whether or not it is.

OpenStudy (kinggeorge):

However, if we're assuming -1 is transcendental, then pi would have to be transcendental for \(e^{i\pi}\) to be transcendental.

OpenStudy (inkyvoyd):

"Sums, products, powers, etc. of the number π and the number e, except for eπ, Gelfond's constant, which is known to be transcendental: π + e, π − e, π·e, π/e, ππ, ee, πe"

OpenStudy (inkyvoyd):

Wait, I reduced the problem incorrectly.

OpenStudy (kinggeorge):

I will have to pose this question in class tomorrow. If we were to somehow call e algebraic, would pi remain transcendental?

OpenStudy (kinggeorge):

The easiest question would be: Given the field \(\mathbb{Q}(e)\), is \(\pi\) algebraic or transcendental over that field?

OpenStudy (kinggeorge):

But if we instead use the field \(\mathbb{E}\) such that every element in \(\mathbb{E}\) is generated by multiplying by e and adding by e, is pi algebraic in that field.

OpenStudy (kinggeorge):

However, for this to be a field, we would need to include all of the rational numbers anyways. So really we would have to use \(\mathbb{Q}(e)\) for our field no matter what.

hero (hero):

Wait, what are we arguing about again?

OpenStudy (inkyvoyd):

If pi in base e is transcendental or algebraic.

OpenStudy (inkyvoyd):

Now, don't you wonder how we got there?